OpenStudy (anonymous):

the hypotenuse of a right triangle has a length of 12 inches and a leg with a length of 8 inches. what is the length of the other leg

OpenStudy (anonymous):

isn't this c square equals a squared plus b squared? I'm not sure. ASNinja will help you he/she knows everything haha

OpenStudy (anonymous):

 \[a^2 + b^2 = c^2\]a = 12 b = ? c = 12

OpenStudy (anonymous):

a^2 + b^2 = c^2 8^2 + b^2 = 12^2 64 + b^2 = 144 B^2 = 80 \[b = \sqrt{80}\]

OpenStudy (anonymous):

simplify the radical

OpenStudy (anonymous):

12*12=144 8*8=64 144-64=80\[\sqrt{80}=8.9\]

OpenStudy (anonymous):

\[4\sqrt{5}\]